OpenStudy (anonymous):

2x +y+z=7 and 3x+2y+4z+19 and 5x +3y+2z+17 solve for x

OpenStudy (texaschic101):

do you prefer elimination or substitution ?

OpenStudy (texaschic101):

oh...and please check your equations again...is there supposed to be equal signs in the last 2 ?

OpenStudy (anonymous):

im sorry yea =19 and =17 and we have been using elimination

OpenStudy (texaschic101):

ok...lets get started 2x + y + z = 7 --->(-4) 3x + 2y + 4z = 19 --------------- -8x - 4y - 4z = -28 --- result of multiplying by -4 3x + 2y + 4z = 19 ---------------add -5x - 2y = - 9 3x + 2y + 4z = 19 5x + 3y + 2z = 17 --->(-2) ---------------- 3x + 2y + 4z = 19 -10x - 6y - 4z = - 34 --- result of multiplying by -2 ---------------add -7x - 3y = - 15 -5x - 2y = -9 --->(3) -7x - 3y = -15 --->(-2) ---------------- -15x - 6y = -27 --- result of multiplying by 3 14x + 6y = 30 --- result of multiplying by -2 ----------------add -x = 3 x = -3 now sub -3 in for x in either equation that you eliminated z in -5x - 2y = -9 -5(-3) - 2y = -9 15 - 2y = -9 -2y = -9 - 15 -2y = - 24 y = 12 now sub our info into any of the original equations 2x + y + z = 7 2(-3) + 12 + z = 7 -6 + 12 + z = 7 6 + z = 7 z = 7 - 6 z = 1 now check our answers 3x + 2y + 4z = 19 3(-3) + 2(12) + 4(1) = 19 -9 + 24 + 4 = 19 -9 + 28 = 19 19 = 19 (correct) solution is : x = -3, y = 12, z = 1 any questions ?

OpenStudy (anonymous):

No...Thank you so much...you are awesome!